PLEASE GIVE ME A VISUAL/GEOMETRICAL UNDERSTANDING OF "DEL" AND HAMILTONIAN OPERATOR"ψ".I AM FACING A PROBLEM IN UNDERSTANDING SCHORINDGER WAVE EQUATION WITHOUT THE MATHEMATICAL UNDERSTANDING OF "DEL".

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Del Operator:
When it is applied to a one dimensional function, it denotes the standard derivative.
When it is applied over multidimensional domain field, then it may denote gradient, divergence or curl of field.
In a three dimensional x, y & z, del is defined as:
\triangledown = \frac{\partial }{\partial x}i + \frac{\partial }{\partial y}j +\frac{\partial }{\partial z}k
Hemiltonian Operator:
Hamiltonian is a operator used in quantum mechanics to denote total energy of system.
Schrodinger Hamiltonian is equal to sum of the operator corresponding to the kinetic & potential energy.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty